2x-3y=16 -3x+2y=-19
Answer by LinnW(1048)

2x-3y=16
-3x+2y=-19
Multiply the top equation by 3 and the bottom equation by 2
6x - 9y = 48
-6x + 4y = -38
Add the two equations
-5y = 10
divide each side by -5
y = -2
substitute 2 for y in 2x-3y=16
2x-3(-2)=16
2x+6 = 16
add -6 to each side
2x = 10
divide each side by2
x = 5
So we have x = 5, and y = -2